• Show that the area bounded by the curve, (x/2)^2 /3 + (y/4)^2/3 = 1 is 3π.
1
Expert's answer
2021-06-20T07:20:31-0400

"\frac{(\frac{x}{2})^2}{3}+\frac{(\frac{y}{4})^2}{3}=1\\
\frac{x^2}{12}+\frac{y^2}{48}=1\\
y=\pm\sqrt{48(1-\frac{x^2}{12})}\\
y=\pm2\sqrt{12-x^2}\\"

To find x limit,

put y=0 in the given equation,

"\frac{x^2}{12}+\frac{0^2}{48}=1\\
x=\pm\sqrt{12}=\pm2\sqrt{3}"

area of the given curve

"=4×\int_a^bydx\\
=4×\int_0^{2\sqrt{3}}2\sqrt{12-x^2}dx\\
=8\int_0^{2\sqrt{3}}\sqrt{(2\sqrt{3})^2-x^2}dx\space (\text{By using integration formula)}\\
=8[\frac{x}{2}\sqrt{12-x^2}+\frac{12}{2}sin^{-1}(\frac{x}{2\sqrt{3}})+c]_0^{2\sqrt{3}}\\
=8[0+6.\frac{\pi}{2}+c-0-0-c]\\
=24\pi" Thus, the area of the given curve is 24π.